OpenStudy (anonymous):

Express in trams of sums and differences of logarithms loga^5x5yz^6

OpenStudy (ranga):

log(P*Q) = log(P) + log(Q)

OpenStudy (anonymous):

explain more plz

OpenStudy (ranga):

log(a^5x5yz^6) = log(a^5) + log(x) + log(5) + log(y) + log(z^6) = 5log(a) + log(x) + log(5) + log(y) + 6log(z)